OpenStudy (spanishrose):

Which ordered pair is a solution of the equation y=x-3? A) (-2,5) B) (-5,2) C) (2,5) D) (5,2)

OpenStudy (agent0smith):

The points are given as (x, y) So plug in each choices x and y values into y=x-3 and see if it's correct.

OpenStudy (spanishrose):

Okay so I think it's 2,5

OpenStudy (spanishrose):

Am I wrong??

OpenStudy (agent0smith):

plug in x=2 and y=5 into y=x-3 and what do you get?

OpenStudy (spanishrose):

2

OpenStudy (agent0smith):

There should be an equals sign in the middle Show all your work.

OpenStudy (spanishrose):

Nevermind, it's not 2

OpenStudy (spanishrose):

2 = 5 + -3 5 + -3 = 2 2 = 2 2 + -2 = 2 + -2 2 + -2 = 0 0 = 2 + -2 2 + -2 = 0 0 = 0 so zero

OpenStudy (spanishrose):

Wait

OpenStudy (spanishrose):

I did my own equation wrong?? I added a plus sign

OpenStudy (agent0smith):

Which point are you doing? Remember the points are (x, y). x first, y second.

OpenStudy (agent0smith):

And you don't solve the equation. 2 = 5 + -3 5 + -3 = 2 2 = 2 Is that true or false? Is 2 equal to 2?

OpenStudy (spanishrose):

True

OpenStudy (agent0smith):

But again I don't know which point you were doing

OpenStudy (spanishrose):

I am so lost omg I don't even know what I'm doing. I'm just going to start over and take your initial advice because I have no idea where I veered off

OpenStudy (agent0smith):

Start with the first point. Plug it into the equation.

OpenStudy (spanishrose):

I think now it's D, 5,2

OpenStudy (agent0smith):

To be sure: plug in x=5, y=2 y=x-3 2 = 5-3 Is that true?

OpenStudy (spanishrose):

Yes

OpenStudy (agent0smith):

Yes.

rebeccaxhawaii (rebeccaxhawaii):

gg